• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 17:34:21 on March16,2012

    Entry number 366735

    keyword=Beam positions adjusted

    Asked MCC to adjust beam positions at IPM1H05A/B. They manage to get it approximately to where we were before: A: 0.5/2.0, B -0.9/0.8